# Predicting Drug-Gene Relations via Analogy Tasks with Word Embeddings

### v1

@@ Introduction
The structure of this paper is as follows. First, we explain related work and detail our approach in this study. In the following sections, we perform experiments for each setting and evaluate the performance using metrics such as top-1 accuracy. Finally, we discuss specific prediction results and then conclude.
@@
In the biomedical field, capturing the relations between biomedical concepts such as drug-drug and protein-protein interaction is an important issue. Word embeddings trained on biomedical domain corpora have been used to predict drug-drug interactions for new drugs [19] and to construct networks of gene-gene interactions [20]. Also, neural networks have been used in a number of studies to predict these relations [21, 22, 23, 24, 25].
Technologies developed in NLP, not limited to word embedding, have been applied to a wide variety of problems in the biomedical domain: Müller et al. [26] have created an online literature search and curation system using an “ontology dictionary” obtained by text mining, and Friedman et al. [27] have used syntactic parsing to extract and structure information about cellular pathways from biological literature. Yeganova et al. [28] normalize synonyms in the biomedical literature by word embedding similarity. Furthermore, Du et al. [29] applied the word embedding algorithm to gene co-expression across the transcriptome to compute vector representations of genes. In recent years, there has also been intensive research on fine-tuning BERT [4] and other pre-trained language models to the biomedical domain [10, 30, 31].
@@ Methods
In this section, based on analogy tasks for drug-gene pairs, we explain analogy tasks in the setting where datasets are divided by year. In this setting, we can separate the drug-gene relations into known or unknown relations based on the chronological order they appear within the PubMed abstracts. By training skip-gram embeddings with this dataset, we can test if embeddings of known relations have the ability to predict unknwon relations. Similar to analogy tasks for drug-gene pairs, we first consider the global setting by year, where all drugs and genes are used in datasets divided by year. Then we consider the pathway-wise setting by year, where drugs and genes are categorized based on pathways in datasets divided by year.
@@
As a first step, based on analogy tasks for drug-gene pairs, we consider analogy tasks in the global setting, where all drugs and genes are used in datasets divided by year. Specifically, drug-gene relations that appeared up to year $y$ are considered known, while those that appeared after year $y$ are considered unknown. We then train skip-gram embeddings using PubMed abstracts published before the specified year, redefine vectors representing known relations, and use these vectors to predict unknown relations. In preparation, we define $y_{d}$ as the year when a drug $d$ first appeared in a PubMed abstract, and $y_{g}$ as the year when a gene $g$ first appeared in a PubMed abstract. In addition, we define $y_{(d,g)}$ as the year when both drug $d$ and gene $g$ first appeared together in a PubMed abstract, and use $y_{(d,g)}$ as a substitute for the year when the relation $(d,g)$ was first identified. By definition, $\max\{y_{d},y_{g}\}\leq y_{(d,g)}$ holds.
@@
Since the vocabulary set of the trained embeddings includes both $d$ and $g$ for $(d,g)\in\mathcal{R}^{y}$, we may discuss this relation $(d,g)$. The relation $(d,g)$ is interpreted as either known by year $y$ if $y_{(d,g)}\leq y$ or unknown by year $y$ if $y<y_{(d,g)}$.
We divide $\mathcal{R}^{y}$ into two subsets based on whether $y_{(d,g)}\leq y$ or $y<y_{(d,g)}$. Let us define two intervals, $L_{y}:=(-\infty,y]$ and $U_{y}:=(y,\infty)$. Considering the set for an interval $I$, consisting of all $(d,g)\in\mathcal{R}^{y}$ such that $y_{(d,g)}\in I$, i.e.,
$\mathcal{R}^{y\mid I}:=\{(d,g)\in\mathcal{R}^{y}\mid y_{(d,g)}\in I\}\subset\mathcal{R}^{y},$ (19)
the set of “known” relations is expressed as $\mathcal{R}^{y\mid L_{y}}$, and the set of “unknown” relations is expressed as $\mathcal{R}^{y\mid U_{y}}$. By definition, $\mathcal{R}^{y\mid L_{y}}\cap\mathcal{R}^{y\mid U_{y}}=\emptyset$ and $\mathcal{R}^{y\mid L_{y}}\cup\mathcal{R}^{y\mid U_{y}}=\mathcal{R}^{y\mid(-\infty,\infty)}=\mathcal{R}^{y}$.
In analogy tasks, we use “known” $\mathcal{R}^{y\mid L_{y}}$ and then predict the target genes $g$ from a drug $d$ for $(d,g)$ in “unknown” $\mathcal{R}^{y\mid U_{y}}$. Using the vector $\mathbf{v}^{y}$, which represents the drug-gene relations and is derived from $\mathcal{R}^{y\mid L_{y}}$, we predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}^{y}$ to $\mathbf{u}_{d}$:
@@
Equation (20) corresponds to Eq. (4). Therefore, similar to the estimator $\hat{\mathbf{v}}$ in Eq. (7), we define an estimator $\hat{\mathbf{v}}^{y}$ for the relation vector $\mathbf{v}^{y}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}^{y\mid L_{y}}$:
$\hat{\mathbf{v}}^{y}:=\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{\left|\mathcal{R}^{y\mid L_{y}}\right|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}(\mathbf{u}_{g}-\mathbf{u}_{d}),$ (21)
where $\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}^{y\mid L_{y}}$. For easy comparison with Eq. (5), Eq. (21) is rewritten as the difference of mean vectors:
$\hat{\mathbf{v}}^{y}=\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{d}\},$ (22)
$\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}^{y\mid L_{y}}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}^{y\mid L_{y}}|}\sum_{(d,g)\in\mathcal{R}^{y\mid L_{y}}}\mathbf{u}_{d}.$ (23)
Given the above, similar to setting G, we perform the analogy tasks in the following two settings.
@@
For the target genes that have drug-gene relations with a drug $d$, only genes whose relations appeared after year $y$ are considered correct. In other words, only new discoveries are counted as correct.
@@
For the target genes that have drug-gene relations with a drug $d$, these genes are considered correct regardless of whether their relations appeared up to or after year $y$. In other words, rediscoveries of known relations are also counted as correct.
See Supplementary Information 1.2.1 for more details on these settings. Note that $y_{d}$, $y_{g}$, and $y_{(d,g)}$ are defined based on the year the drugs, genes, and drug-gene relations appeared in PubMed abstracts. Thus, they do not fully correspond to their actual years of discovery, and we only consider the analogy tasks in these hypothetical settings.
@@
As a next step, based on the analogy tasks in settings P1, P2, Y1, and Y2, we consider more detailed analogy tasks. To do this, we consider the analogy tasks in the pathway-wise setting by year, where drugs and genes are categorized based on pathways in datasets divided by year.
Similar to settings P1, P2, Y1, and Y2, we perform the analogy tasks in the following four settings.
@@
For the target genes that have drug-gene relations with a drug $d$, only genes categorized in the same pathway $p$ as $d$, and whose relations appeared after year $y$, are considered correct.
@@
For the target genes that have drug-gene relations with a drug $d$, only genes categorized in the same pathway $p$ as the drug $d$ are considered correct, regardless of whether their relations appeared up to or after year $y$.
Setting P2Y1.
For the target genes that have drug-gene relations with a drug $d$, only genes whose relations appeared after year $y$ are considered correct, regardless of whether they are categorized in the same pathway $p$ as the drug $d$ or not.
@@
For the target genes that have drug-gene relations with a drug $d$, these genes are considered correct regardless of whether they are categorized in the same pathway or not, and whether their relations appeared up to or after year $y$.
See Supplementary Information 1.2.2 for details on the analogy tasks and these settings.
@@
Drug-gene relations
@@
Baseline
To evaluate the performance of predicting target genes by adding the relation vector to a drug embedding, we compare it to a simple baseline method. As a baseline, we randomly select one gene from the search space for each query. We refer to this baseline as the random baseline. In the random baseline, we repeated the experiments 10 times and used the average score as the final result.
@@
As evaluation metrics, we use the top-$k$ accuracy explained in each setting, especially the top-1 and top-10 accuracies. We also use Mean Reciprocal Rank (MRR) as another evaluation metric. MRR is a statistical measure to evaluate search performance, expressed as the sum of the reciprocals of the ranks at which the predicted results first appear in the answer set.
When evaluating performance using top-1 accuracy, top-10 accuracy, and MRR, the embeddings of all genes $\mathcal{G}$ (or $\mathcal{G}^{y}$ in settings by year) are centered using the mean embedding $\mathrm{E}_{\mathcal{G}}\{\mathbf{u}_{g}\}$ in Eq. (6) (or $\mathrm{E}_{\mathcal{G}^{y}}\{\mathbf{u}_{g}\}$). Since cosine similarity is affected by the origin, centering mitigates this effect. In the case of setting G, for example, we actually calculate the value of $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}}-\mathrm{E}_{\mathcal{G}}\{\mathbf{u}_{g}\},\mathbf{u}_{g}-\mathrm{E}_{\mathcal{G}}\{\mathbf{u}_{g}\})$ instead of $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}},\mathbf{u}_{g})$.
@@ Results
Embeddings Setting Method Top1 Top10 MRR
@@
Gene prediction performance in settings G, P1, and P2.
Table 3 shows the results of experiments with BioConceptVec and our skip-gram embeddings in settings G, P1, and P2. In setting G, the prediction by simply adding the global relation vector $\hat{\mathbf{v}}$ resulted in a top-1 accuracy of about 0.3, a top-10 accuracy of over 0.6, and an MRR also over 0.4. These results show that the embeddings have the ability to interpret the drug-gene relations. Similar to the basic analogy task, these relations were not explicitly provided during the training of the embeddings. Previous research [37] has shown that skip-gram achieves a top-1 accuracy of 0.56 on the Microsoft Research Syntactic Analogies Dataset [38]. Given this, BioConceptVec and our skip-gram embeddings in the analogy tasks for drug-gene pairs show comparable performance to skip-gram in the basic analogy tasks.
@@
In contrast, the random baseline in settings G, P1, and P2 showed nearly zero scores for all evaluation metrics, suggesting that random prediction is nearly impossible in these analogy tasks. This also confirms the superior performance of prediction by adding relation vectors.
In addition, our skip-gram outperformed BioConceptVec except for the top-1 accuracy in setting G. This is probably because BioConceptVec has a dimensionality of 100, while our skip-gram has a dimensionality of 300. Note that, as shown in Table 2, the settings of the analogy tasks are not identical for these two embeddings.
@@
Figures 3(a) and 3(b) show the results of experiments with our skip-gram embeddings in settings Y1 and Y2. BioConceptVec embeddings are not appropriate for the settings by year, since they are pretrained on the entire dataset without year-specific divisions.
@@
In setting Y1, since only genes whose relations appeared after year $y$ are included in the answer set, the tasks are more challenging. Therefore, the evaluation metric scores are lower compared to those in setting Y2, where genes whose relations appeared up to year $y$ are also included in the answer set. Yet, in setting Y1, the top-1 accuracy from $y=1985$ to $y=2015$ is about 0.1, the top-10 accuracy is over 0.3, and the MRR is about 0.2. Although in setting Y1 the relations to be predicted appeared after year $y$ and are not used to calculate $\hat{\mathbf{v}}^{y}$, these results show that adding $\hat{\mathbf{v}}^{y}$ to $\mathbf{u}_{d}$ can predict these relations. Setting Y2 is easier than setting Y1, and using the relation vectors results in high evaluation metric scores. For example, from $y=1985$ to $y=2015$, the top-1 accuracy reaches about 0.3, and the top-10 accuracy is about 0.7.
@@
In setting Y1, the number of queries increases initially and then decreases over time, while in setting Y2, the number of queries increases steadily from older to more recent years. In addition, the size of the search space is small for older years (e.g., $|\mathcal{G}^{1970}|=353$) and increases for later years (e.g., $|\mathcal{G}^{2020}|=47509$). As a result, in both settings Y1 and Y2, the datasets for older years such as $y=1970$, $1975$, and $1980$ have fewer queries and smaller search spaces, resulting in unusually high evaluation metric scores. In setting Y1, the datasets for later years have fewer queries and larger search spaces, which tends to produce lower evaluation metric scores. In both settings, the evaluation metric scores for the random baseline are close to zero. For more detailed results for settings Y1 and Y2, see Table S8 in Supplementary Information 2.2.1.
Gene prediction performance in settings P1Y1, P1Y2, P2Y1 and P2Y2.
@@
Figures 3(c) and 3(d) show the results of experiments with our skip-gram embeddings in settings P1Y1 and P1Y2. The results in settings P1Y1 and P1Y2 not only follow similar trends to those in settings Y1 and Y2 but also consistently show higher evaluation metric scores. The analogy tasks in settings P1Y1 and P1Y2 use detailed pathway information. Therefore, predicting the relations that appeared after year $y$ is easier in settings P1Y1 and P1Y2 than in settings Y1 and Y2. In addition, the results for the older years show extremely high evaluation metric scores in settings P1Y1 and P1Y2, similar to those in settings Y1 and Y2. The evaluation metric scores of the random baseline are close to zero.
@@
For the results in settings P2Y1 and P2Y2, see Fig. S2 in Supplementary Information 2.2.2. They are very similar to those in settings P1Y1 and P1Y2. For detailed results for settings P1Y1, P1Y2, P2Y1, and P2Y2, see Table S10 in Supplementary Information 2.2.2.
Discussion
@@
In this section, we investigate the predictions made by adding the relation vector in settings G, P1, and P2. For this purpose, we focus on the results using genes and drugs categorized in the ErbB signaling pathway as shown in Table 1 and Fig. 2. Furthermore, since our skip-gram embeddings showed better performance compared to the BioConceptVec skip-gram embeddings (Table 3), we adopted our embeddings for this analysis. $\mathcal{D}_{p}$, $\mathcal{G}_{p}$, and $\mathcal{R}_{p}$ of our embeddings are identical to those of the BioConceptVec embeddings in Table 1.
@@
In Table 4, we list the predicted target genes for drugs that are categorized in the ErbB signaling pathway and whose answer set size in setting G is two or more. For all drugs meeting these criteria, at least one answer gene was included in the top 10 predicted target genes for either of the settings. In the following, we show that genes that were ranked high in the predictions, even though they were not in the answer set, can be interpreted biologically.
@@
First, the target genes of Bosutinib are ABL1 and SRC, which are both known to be non-receptor tyrosine kinases (non-RTKs) [39]. Within the top 10 predicted target genes, TXK [40] and JAK2 [41] can also be categorized as non-RTKs, sharing some properties with ABL1 and SRC. This implies that the structural and biochemical similarities of these genes may have been reflected in their high-dimensional representations.

### v2

@@ Introduction
The structure of this paper is as follows. First, we explain related work and detail our approach in this study. In the following sections, we perform experiments for each setting and evaluate the performance using metrics such as top-1 accuracy. Finally, we discuss the prediction results and then conclude.
@@
In the biomedical field, capturing the relations between biomedical concepts, such as drug-drug and protein-protein interactions, is an important issue. Word embeddings trained on biomedical domain corpora have been used to predict drug-drug interactions for new drugs [19] and to construct networks of gene-gene interactions [20].
While word embeddings have shown promise in various biomedical applications, their potential for predicting drug-gene relations remains relatively unexplored. To the best of our knowledge, there is limited research specifically focused on utilizing analogy computation of word embeddings to predict drug-gene relations in the biological domain. Our study aims to address this gap by investigating the application of analogy computation of word embeddings to drug-gene relationship prediction.
In addition, neural networks have been used in a number of studies to predict these relations [21, 22, 23, 24, 25]. However, it is important to note that these relation extraction approaches differ from our proposed task. While they are useful for identifying explicitly stated relationships in text, our focus is on leveraging the latent semantic information captured by word embeddings to predict potential drug-gene relations, even when such relationships are not explicitly mentioned in the literature.
The technologies developed in NLP, not limited to word embeddings, have been applied to a wide variety of problems in the biomedical domain, demonstrating their utility in this field. Müller et al. [26] have created an online literature search and curation system using an “ontology dictionary” obtained by text mining, and Friedman et al. [27] have used syntactic parsing to extract and structure information about cellular pathways from biological literature. Yeganova et al. [28] normalize synonyms in the biomedical literature by word embedding similarity. Furthermore, Du et al. [29] applied the word embedding algorithm to gene co-expression across the transcriptome to compute vector representations of genes. In recent years, there has also been intensive research on fine-tuning BERT [4] and other pre-trained language models to the biomedical domain [10, 30, 31].
@@ Methods
In this section, based on analogy tasks for drug-gene pairs, we explain analogy tasks in the setting where datasets are divided by year. To do so, we first use embeddings trained on PubMed abstracts up to year $y$ and consider analogy tasks in a global setting such as setting G. Next, we separate the drug-gene relations into “known” or “unknown” based on their chronological appearance in the PubMed abstracts. Using the embeddings from the datasets divided by year, we test whether embeddings of “known” relations have the ability to predict “unknwon” relations. We then consider the pathway-wise settings by year, where drugs and genes are categorized based on pathways in the datasets divided by year. In these settings, we use the year-specific embeddings and consider analogy tasks in settings such as P1 and P2. Finally, we evaluate whether “unknown” relations can be predicted by “known” relations.
@@
First, using embeddings trained on PubMed abstracts up to year $y$, we consider analogy tasks in a global setting. In preparation, we define $y_{d}$ as the year when a drug $d$ first appeared in a PubMed abstract and $y_{g}$ as the year when a gene $g$ first appeared in a PubMed abstract.
@@
Similar to the global setting, for drug-gene pairs $(d,g)\in\mathcal{R}^{y}$, we consider the analogy tasks for predicting the target genes $g$ from a drug $d$. To solve these analogy tasks, we use the relation vector $\mathbf{v}^{y}$. We predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}^{y}$ to $\mathbf{u}_{d}$:
@@
Equation (19) corresponds to Eq. (4). Therefore, similar to the estimator $\hat{\mathbf{v}}$ in Eq. (7), we define an estimator $\hat{\mathbf{v}}^{y}$ for the relation vector $\mathbf{v}^{y}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}^{y}$:
$\hat{\mathbf{v}}^{y}:=\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{\left|\mathcal{R}^{y}\right|}\sum_{(d,g)\in\mathcal{R}^{y}}(\mathbf{u}_{g}-\mathbf{u}_{d}),$ (20)
where $\mathrm{E}_{\mathcal{R}^{y}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}^{y}$. For easy comparison with Eq. (5), Eq. (20) is rewritten as the difference of mean vectors:
$\hat{\mathbf{v}}^{y}=\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{d}\},$ (21)
$\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}^{y}|}\sum_{(d,g)\in\mathcal{R}^{y}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}^{y}|}\sum_{(d,g)\in\mathcal{R}^{y}}\mathbf{u}_{d}.$ (22)
Given the above, similar to setting G, we perform the analogy tasks in the following setting.
@@
Using embeddings trained on PubMed abstracts up to year $y$, we consider analogy tasks in setting G. Thus, if $y$ is the most recent, it simply corresponds to setting G.
See Supplementary Information 1.2.1 for more details on the setting.
Global setting to predict unknown relations by year
Based on analogy tasks in the global setting by year, we consider analogy tasks to predict unknown relations using known relations. Specifically, drug-gene relations that appeared up to year $y$ are considered known, while those that appeared after year $y$ are considered unknown. We then use embeddings trained on PubMed abstracts up to year $y$, redefine vectors representing known relations, and use these vectors to predict unknown relations. In preparation, we define $y_{(d,g)}$ as the year when both drug $d$ and gene $g$ first appeared together in a PubMed abstract. We consider $y_{(d,g)}$ as a substitute for the year when the relation $(d,g)$ was first identified. By definition, $\max\{y_{d},y_{g}\}\leq y_{(d,g)}$ holds. The relation $(d,g)$ is interpreted as either known by year $y$ if $y_{(d,g)}\leq y$ or unknown by year $y$ if $y<y_{(d,g)}$.
We define two subsets of $\mathcal{R}^{y}$ based on whether $y_{(d,g)}\leq y$ or $y<y_{(d,g)}$. To do this, we define two intervals, $L_{y}:=(-\infty,y]$ and $U_{y}:=(y,\infty)$. Using $L_{y}$ and $U_{y}$, we define the subsets $\mathcal{R}^{y\mid L_{y}}$ and $\mathcal{R}^{y\mid U_{y}}$ of $\mathcal{R}^{y}$ as follows:
$\mathcal{R}^{y\mid L_{y}}:=\{(d,g)\in\mathcal{R}^{y}\mid y_{(d,g)}\in L_{y}\},\,\mathcal{R}^{y\mid U_{y}}:=\{(d,g)\in\mathcal{R}^{y}\mid y_{(d,g)}\in U_{y}\}\subset\mathcal{R}^{y}.$ (23)
The set of “known” relations is expressed as $\mathcal{R}^{y\mid L_{y}}$, and the set of “unknown” relations is expressed as $\mathcal{R}^{y\mid U_{y}}$. By definition, $\mathcal{R}^{y\mid L_{y}}\cap\mathcal{R}^{y\mid U_{y}}=\emptyset$ and $\mathcal{R}^{y\mid L_{y}}\cup\mathcal{R}^{y\mid U_{y}}\subset\mathcal{R}^{y}$.
In analogy tasks, we use “known” $\mathcal{R}^{y\mid L_{y}}$ and then predict the target genes $g$ from a drug $d$ for $(d,g)$ in “unknown” $\mathcal{R}^{y\mid U_{y}}$. Using the relation vector $\mathbf{v}^{y\mid L_{y}}$, which represents the drug-gene relations in $\mathcal{R}^{y\mid L_{y}}$, we predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}^{y}$ to $\mathbf{u}_{d}$:
$\mathbf{u}_{d}+\mathbf{v}^{y\mid L_{y}}\approx\mathbf{u}_{g}.$ (24)
Equation (24) corresponds to Eq. (19). We define the estimator $\hat{\mathbf{v}}^{y\mid L_{y}}$ for the relation vector $\mathbf{v}^{y\mid L_{y}}$ by using $\mathcal{R}^{y\mid L_{y}}$ instead of $\mathcal{R}^{y}$ in the estimator $\hat{\mathbf{v}}^{y}$ in Eq. (20). Given the above, we perform the analogy tasks in the following setting.
@@
Using embeddings trained on PubMed abstracts up to year $y$, we consider analogy tasks with a reduced answer set. For the target genes that have drug-gene relations with a drug $d$, only genes whose relations appeared after year $y$ are considered correct. In other words, only new discoveries are counted as correct.
See Supplementary Information 1.2.2 for more details on the setting. Note that $y_{d}$, $y_{g}$, and $y_{(d,g)}$ are defined based on the year the drugs, genes, and drug-gene relations appeared in PubMed abstracts. Thus, they do not fully correspond to their actual years of discovery, and we only consider the analogy tasks in these hypothetical settings.
@@
As a next step, based on the analogy tasks in settings P1, P2, and Y1, we consider the analogy tasks in the pathway-wise setting by year, where drugs and genes are categorized based on pathways in datasets divided by year. To do this, we perform the analogy tasks in the following two settings.
@@
Using embeddings trained on PubMed abstracts up to year $y$, we consider analogy tasks in setting P1. Thus, if $y$ is the most recent, it corresponds to setting P1.
Setting P2Y1.
Using embeddings trained on PubMed abstracts up to year $y$, we consider analogy tasks in setting P2. Thus, if $y$ is the most recent, it corresponds to setting P2.
See Supplementary Information 1.2.3 for details on the analogy tasks and these settings.
Pathway-wise setting to predict unknown relations by year
Furthermore, based on the analogy tasks in settings P1, P2, and Y2, we consider analogy tasks to predict unknown relations using known relations in the pathway-wise setting by year. To do this, we perform the analogy tasks in the following two settings.
@@
Using embeddings trained on PubMed abstracts up to year $y$, we consider analogy tasks in setting P1 with a reduced answer set. For the target genes that have drug-gene relations with a drug $d$, only genes categorized in the same pathway $p$ as $d$, and whose relations appeared after year $y$, are considered correct.
@@
Using embeddings trained on PubMed abstracts up to year $y$, we consider analogy tasks in setting P2 with a reduced answer set. For the target genes that have drug-gene relations with a drug $d$, only genes whose relations appeared after year $y$ are considered correct, regardless of whether they are categorized in the same pathway $p$ as the drug $d$ or not.
See Supplementary Information 1.2.4 for details on the analogy tasks and these settings.
@@
Drug-gene relations.
@@
Baselines
To evaluate the performance of predicting target genes by adding the relation vector to a drug embedding, we compare it to baseline methods. As a simple baseline, we randomly select one gene from the search space for each query. We refer to this baseline as the random baseline. In the random baseline, we repeated the experiments 10 times and used the average score as the final result. In addition to this, as strong generative model baselines, we used the Generative Pre-trained Transformer (GPT) [37] series to predict the top 10 target genes from a query drug by zero-shot. We used the GPT-3.5 [38], GPT-4 [39], and GPT-4o (https://platform.openai.com/docs/models/gpt-4o) models with the temperature hyperparameter set to 00. See Supplementary Information 1.6 for more details on the models and prompt template for the predictions.
@@
As evaluation metrics, we use the top-$k$ accuracy explained in each setting, especially the top-1 and top-10 accuracies. We also use Mean Reciprocal Rank (MRR) as another evaluation metric. MRR is a statistical measure to evaluate search performance, expressed as the sum of the reciprocals of the ranks at which the predicted results first appear in the answer set. Note that since MRR cannot be calculated for the GPT models due to the limited number of prediction candidates output by GPT, only top-1 and top-10 accuracies are calculated.
When evaluating performance using top-1 accuracy, top-10 accuracy, and MRR, the embeddings of all genes $\mathcal{G}$ (or $\mathcal{G}^{y}$ in settings by year) are centered using the mean embedding $\mathrm{E}_{\mathcal{G}}\{\mathbf{u}_{g}\}$ in Eq. (6) (or $\mathrm{E}_{\mathcal{G}^{y}}\{\mathbf{u}_{g}\}$). Since cosine similarity is affected by the origin, centering mitigates this effect. In the case of setting G, for example, we actually calculate the value of $\cos(\mathbf{u}_{d}-\mathrm{E}_{\mathcal{G}}\{\mathbf{u}_{g}\}+\hat{\mathbf{v}},\mathbf{u}_{g}-\mathrm{E}_{\mathcal{G}}\{\mathbf{u}_{g}\})$ instead of $\cos(\mathbf{u}_{d}+\hat{\mathbf{v}},\mathbf{u}_{g})$.
@@ Results
Model Setting Method Top1 Top10 MRR
@@
GPT-3.5 Best of all zero-shot 0.653 0.821
GPT-4 Best of all zero-shot 0.710 0.832
GPT-4o Best of all zero-shot 0.760 0.887
@@
Gene prediction performance in settings G, P1, and P2
Table 3 shows the results of experiments with BioConceptVec and our skip-gram embeddings in settings G, P1, and P2. For GPT-3.5, GPT-4, and GPT-4o, only the best results are shown; see Supplementary Information 2.1.3 for detailed results. In setting G, the prediction by simply adding the global relation vector $\hat{\mathbf{v}}$ resulted in a top-1 accuracy of about 0.3, a top-10 accuracy of over 0.6, and an MRR also over 0.4. These results show that the embeddings have the ability to interpret the drug-gene relations. As with the basic analogy task, these relations were not explicitly provided during the training of the embeddings.
@@
The random baseline in settings G, P1, and P2 showed nearly zero scores for all evaluation metrics, suggesting that random prediction is nearly impossible in these analogy tasks. This also confirms the superior performance of prediction by adding relation vectors. Among the GPT series, GPT-4o performed the best, followed by GPT-4, which outperformed GPT-3.5. In top-1 accuracy, the GPT series performed better than predictions made by adding relation vectors. However, in top-10 accuracy, the difference becomes smaller. Notably, predictions using $\mathbf{v}_{p}$ with our skip-gram embeddings in settings P1 and P2 performed comparably to GPT-4o and outperformed both GPT-3.5 and GPT-4.
In addition, our skip-gram outperformed BioConceptVec except for the top-1 accuracy in setting G. This is probably because BioConceptVec has a dimensionality of 100 and the size of the search space is $|\mathcal{G}|=117282$, while our skip-gram has a dimensionality of 300 and the size of the search space is $|\mathcal{G}|=28284$, as shown in Table 2.
@@
Figures 3(a) and 3(b) show the results of experiments in settings Y1 and Y2 with our skip-gram embeddings trained on the year-specific divisions up to year $y$. BioConceptVec embeddings are not appropriate for the settings by year, since they are pretrained on the entire dataset without year-specific divisions.
Setting Y1 is easier than setting Y2, and results in high evaluation metric scores. For example, from $y=1980$ to $y=2023$, the top-1 accuracy reaches about 0.3, and the top-10 accuracy is about 0.6. In setting Y1, as in setting G, all relations contained in $\cal{R}$ are considered correct, so all the scores for $y=2023$ are identical to those of setting G. For example, a top-10 accuracy of 0.686 is confirmed in both Fig. 3(a) and Table 3.
@@
In setting Y2, since only genes whose relations appeared after year $y$ are included in the answer set, the tasks are more challenging. Therefore, the evaluation metric scores are lower compared to those in setting Y1, where genes whose relations appeared up to year $y$ are also included in the answer set. Yet, in setting Y2, the top-1 accuracy from $y=1985$ to $y=2015$ is about 0.1, the top-10 accuracy is over 0.3, and the MRR is about 0.2. Although in setting Y2 the relations to be predicted appeared after year $y$ and are not used to calculate $\hat{\mathbf{v}}^{y\mid L_{y}}$, these results show that adding $\hat{\mathbf{v}}^{y\mid L_{y}}$ to $\mathbf{u}_{d}$ can predict these relations.
@@
In setting Y1, the number of queries increases steadily from older to more recent years, while in setting Y2, the number of queries increases initially and then decreases over time. In addition, the size of the search space is small for older years (e.g., $|\mathcal{G}^{1975}|=725$) and increases for later years (e.g., $|\mathcal{G}^{2020}|=47509$). As a result, in both settings Y1 and Y2, the datasets for older years such as $y=1975$, and $1980$ have fewer queries and smaller search spaces, resulting in unusually high evaluation metric scores. In setting Y2, the datasets for later years have fewer queries and larger search spaces, which tends to produce lower evaluation metric scores. In both settings, the evaluation metric scores for the random baseline are close to zero. For more detailed results for settings Y1 and Y2, see Table S12 in Supplementary Information 2.2.1.
Gene prediction performance in settings P1Y1, P2Y1, P1Y2, and P2Y2
Figures 3(c) and 3(d) show the results of experiments with our skip-gram embeddings in settings P1Y1 and P1Y2. The results in settings P1Y1 and P1Y2 not only follow similar trends to those in settings Y1 and Y2 but also consistently show higher evaluation metric scores. From the definition of setting P1Y1, the scores for $y=2023$ are identical to those of setting P1. For example, a top-10 accuracy of 0.862 is confirmed in both Fig. 3(c) and Table 3.
@@
The analogy tasks in settings P1Y1 and P1Y2 use detailed pathway information. Therefore, predicting the relations that appeared after year $y$ is easier in settings P1Y1 and P1Y2 than in settings Y1 and Y2. In addition, the results for the older years show high evaluation metric scores in settings P1Y1 and P1Y2, similar to those in settings Y1 and Y2. The evaluation metric scores of the random baseline are close to zero.
@@
For the results in settings P2Y1 and P2Y2, see Fig. S2 in Supplementary Information 2.2.2. They are very similar to those in settings P1Y1 and P1Y2. For detailed results for settings P1Y1, P2Y1, P1Y2, and P2Y2, see Table S15 in Supplementary Information 2.2.2.
Biological Insights from Predicting Drug-Gene Relations
@@
In this section, we investigate the biological insights obtained through analogy tasks by adding the relation vector in settings G, P1, and P2. For this purpose, we focus on genes and drugs categorized in the ErbB signaling pathway as shown in Table 1 and Fig. 2. Furthermore, since our skip-gram embeddings showed better performance compared to the BioConceptVec skip-gram embeddings (Table 3), we adopted our embeddings for this analysis. $\mathcal{D}_{p}$, $\mathcal{G}_{p}$, and $\mathcal{R}_{p}$ of our embeddings are identical to those of the BioConceptVec embeddings in Table 1.
@@
In Table 4, we have listed the predicted target genes for drugs that are categorized in the ErbB signaling pathway and whose answer set size in setting G is two or more. For all drugs meeting these criteria, at least one answer gene was included in the top 10 predicted target genes for any of the settings. Therefore, it can be said that the prediction of drug-gene relations through analogy tasks functioned appropriately. Additionally, we provide several examples demonstrating that some genes ranked high in the predictions, even though they were not in the answer set, can still be interpreted biologically.
@@
First, the target genes of Bosutinib are ABL1 and SRC, which are both known to be non-receptor tyrosine kinases (non-RTKs) [40]. Among the top 10 predicted target genes, TXK [41] and JAK2 [42] can also be categorized as non-RTKs, sharing several properties with ABL1 and SRC. This implies that the structural and biochemical similarities of these genes may have been reflected in their high-dimensional representations.